-10 micro Columb,40 micro column and q are the charges on three identical spherical conductors P,Q and R respectively.Now P and Q attract each other with a force F when they are separated by a distance d.Now P and Q are made in contact with each other and then separated.Again Q and R are touched and they are separated by a distance d.The repulsive force between Q and R is 4F.Then the charge q is

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

As given.-F=\dfrac{k(-10)(40)}{d^2}
When P and Q are made in contact, charge on each of them is\dfrac{-10+40}{2}C=+15C
Now when Q and R are touched, the charge on each of them is\dfrac{+15+q}{2}C=
When these are separated by distanced, repulsive force acting on them is

\dfrac{k(15+q)^2}{4d^2}=4F
\implies q=+65COR-95C